# Game-Theoretic Multiagent Reinforcement Learning

### v4

@@ 8.3 Solutions by Two-Timescale Analysis
Two-timescale stochastic analysis is a set of tools certifying that, in a system with two coupled stochastic processes that evolve at different speeds, if the fast process converges to a unique limit point for any particular fixed value of the slow process, we can, quantitatively, analyse the asymptotic behaviour of the algorithm as if the fast process is always fully calibrated to the current value of the slow process [56]. As a direct application, [318, 319] noted that independent Q-learners with agent-dependent learning rates could break the symmetry that leads to the non-convergent limited cycles; as a result, they can converge almost surely to the NE in two-player collaboration games, two-player zero-sum games, and multi-player matching pennies. Similarly, [436] introduced a two-timescale update rule that ensures the training dynamics reach a stationary local NE in general-sum SGs if the critic learns faster than the actor. Later, [425] proposed a distributed actor-critic algorithm that enjoys provable convergence in solving static potential games with continuous actions. Similarly, [18] developed a two-timescale variant of Q-learning that is guaranteed to converge to an equilibrium in SGs with weakly acyclic characteristics, which generalises potential games. Other applications include developing two-timescale update rules for training GANs [244] and developing a two-timescale algorithm with guaranteed asymptotic convergence to the Stackelberg equilibrium in general-sum Stackelberg games.
@@ 8.4 Solutions by Policy-Based Methods
where $\nabla_{\mathbf{w}_{i},\mathbf{w}_{j}}^{2}\ell_{k}$ is the $(d_{i}\times d_{j})$ block of $2$nd-order derivatives. A differentiable NE for the game is $\bm{w}^{*}$ if $\bm{\xi}(\mathbf{w}^{*})=0$ and $\nabla_{\mathbf{w}_{i}}^{2}\ell_{i}\succ 0,\ \forall i\in\{1,...,N\}$; furthermore, this result is a local NE if $\det\mathbf{H}(\mathbf{w}^{*})\neq 0$.
@@
Figure 9.1: Relations of mean-field learning algorithms in games with large $N$.
@@ 9.1 Non-cooperative Mean-Field Game
The challenge of the above problem lies in the coupled relationship between $T$ and $\tau^{i}$; that is, in order to compute $T$, we need to know $\tau^{i}$, which is based on $T$ itself. Therefore, solving the time $T$ is essentially equivalent to finding the fixed point, if it exists, of the stochastic process that generates $T$. In fact, $T$ can be effectively computed through a two-step iterative process, and we denote as $\Gamma^{1}$ and $\Gamma^{2}$. At $\Gamma^{1}$, given the current [^3] At time step 00, it can be a random guess. Since the fixed point exists, the final convergence result is irrelevant to the initial guess. value of $T$, each agent solves their optimal arrival time $\tau^{i}$ by minimising their cost $R^{i}({t,T,\tilde{\tau}^{i}})$. At $\Gamma^{2}$, agents calibrate the new estimate of $T$ based on all $\tau^{i}$ values that were computed in $\Gamma^{1}$. $\Gamma^{1}$ and $\Gamma^{2}$ continue iterating until $T$ converges to a fixed point, i.e., $\Gamma^{2}\circ\Gamma^{1}(T^{*})=T^{*}$. The key insight is that the interaction with other agents is captured simply by the mean-field quantity. Since the meeting starts only when $90\%$ of the people arrive, if one considers a continuum of players with $N\rightarrow+\infty$, $T$ becomes the $90th$ quantile of a distribution, and each agent can easily find the best response. This result contrasts to the cases of a finite number of players, in which the ordered statistic is intractable, especially when $N$ is large (but still finite).
@@
[…]
@@ 9.2 Cooperative Mean-Field Control
Another way to address the time inconsistency in MFCs is to consider an enlarged state-action space [312, 430, 432, 431, 145, 214]. This technique is also called “lift up”, which essentially means to lift up the state space and the action space into their corresponding probability measure spaces in which dynamic programming principles hold. For example, [214, 396] proposed to lift the finite state-action space $\mathbb{S}$ and $\mathbb{A}$ to a compact state-action space embedded in Euclidean space denoted by $\mathcal{C}:=\Delta(\mathbb{S})\times\mathcal{H}$ and :=$\mathcal{H}:=\big{\{}h:\mathbb{S}\rightarrow\Delta({\mathbb{A}})\big{\}}$, and the optimal Q-function associated with the MFC problem in Eq. (9.4) is
@@ 9.3 Mean-Field MARL
Chapter 10 Future Directions of Interest

### v5

@@ 8.3 Solutions by Two-Timescale Analysis
Two-timescale stochastic analysis is a set of tools certifying that, in a system with two coupled stochastic processes that evolve at different speeds, if the fast process converges to a unique limit point for any particular fixed value of the slow process, we can, quantitatively, analyse the asymptotic behaviour of the algorithm as if the fast process is always fully calibrated to the current value of the slow process [56]. As a direct application, [318, 319] noted that independent Q-learners with agent-dependent learning rates could break the symmetry that leads to the non-convergent limit cycles; as a result, they can converge almost surely to the NE in two-player collaboration games, two-player zero-sum games, and multi-player matching pennies. Similarly, [436] introduced a two-timescale update rule that ensures the training dynamics reach a stationary local NE in general-sum SGs if the critic learns faster than the actor. Later, [425] proposed a distributed actor-critic algorithm that enjoys provable convergence in solving static potential games with continuous actions. Similarly, [18] developed a two-timescale variant of Q-learning that is guaranteed to converge to an equilibrium in SGs with weakly acyclic characteristics, which generalises potential games. Other applications include developing two-timescale update rules for training GANs [244] and developing a two-timescale algorithm with guaranteed asymptotic convergence to the Stackelberg equilibrium in general-sum Stackelberg games.
@@ 8.4 Solutions by Policy-Based Methods
where $\nabla_{\mathbf{w}_{i},\mathbf{w}_{j}}^{2}\ell_{k}$ is the $(d_{i}\times d_{j})$ block of 22nd-order derivatives. A differentiable NE for the game is $\bm{w}^{*}$ if $\bm{\xi}(\mathbf{w}^{*})=0$ and $\nabla_{\mathbf{w}_{i}}^{2}\ell_{i}\succ 0,\ \forall i\in\{1,...,N\}$; furthermore, this result is a local NE if $\det\mathbf{H}(\mathbf{w}^{*})\neq 0$.
@@
Figure 9.1: Relations of mean-field learning algorithms in games with large NN.
@@ 9.1 Non-cooperative Mean-Field Game
The challenge of the above problem lies in the coupled relationship between $T$ and $\tau^{i}$; that is, in order to compute $T$, we need to know $\tau^{i}$, which is based on $T$ itself. Therefore, solving the time $T$ is essentially equivalent to finding the fixed point, if it exists, of the stochastic process that generates $T$. In fact, $T$ can be effectively computed through a two-step iterative process, and we denote as $\Gamma^{1}$ and $\Gamma^{2}$. At $\Gamma^{1}$, given the current [^3] At time step 0, it can be a random guess. Since the fixed point exists, the final convergence result is irrelevant to the initial guess. value of $T$, each agent solves their optimal arrival time $\tau^{i}$ by minimising their cost $R^{i}({t,T,\tilde{\tau}^{i}})$. At $\Gamma^{2}$, agents calibrate the new estimate of $T$ based on all $\tau^{i}$ values that were computed in $\Gamma^{1}$. $\Gamma^{1}$ and $\Gamma^{2}$ continue iterating until $T$ converges to a fixed point, i.e., $\Gamma^{2}\circ\Gamma^{1}(T^{*})=T^{*}$. The key insight is that the interaction with other agents is captured simply by the mean-field quantity. Since the meeting starts only when $90\%$ of the people arrive, if one considers a continuum of players with $N\rightarrow+\infty$, $T$ becomes the $90th$ quantile of a distribution, and each agent can easily find the best response. This result contrasts to the cases of a finite number of players, in which the ordered statistic is intractable, especially when $N$ is large (but still finite).
@@
$\text{ subject to }s_{t+1}$ $\sim P\left(s_{t},a_{t},L_{t}\right),a_{t}\sim\pi_{t}\left(s_{t}\right).$ (9.2)
@@ 9.2 Cooperative Mean-Field Control
Another way to address the time inconsistency in MFCs is to consider an enlarged state-action space [312, 430, 432, 431, 145, 214]. This technique is also called “lift up”, which essentially means to lift up the state space and the action space into their corresponding probability measure spaces in which dynamic programming principles hold. For example, [214, 396] proposed to lift the finite state-action space $\mathbb{S}$ and $\mathbb{A}$ to a compact state-action space embedded in Euclidean space denoted by $\mathcal{C}:=\Delta(\mathbb{S})\times\mathcal{H}$ and $\mathcal{H}:=\big{\{}h:\mathbb{S}\rightarrow\Delta({\mathbb{A}})\big{\}}$, and the optimal Q-function associated with the MFC problem in Eq. (9.4) is
@@ 9.3 Mean-Field MARL
Chapter 10 Future Directions